50 YEARS OF “MATEMATIČKI VESNIK”

By the fall of 1948 the Managing Board of the Society of Mathematicians and Physicists of Serbia decided to start publishing a scientific journal named VESNIK DRUŠTVA MATEMATIČARA I FIZIČARA NR SRBIJE (“Bulletin of the Society of Mathematicians and Physicists of Serbia”). Jovan Karamata was the first editorin-chief, the Editorial Board consisted of Pavle Savić, Dragoljub K. Jovanović, Miloš Radojčić and Dobrivoje Mihajlović, while Ivan Atanasijević was the technical editor.

The first issue of the journal was published in the beginning of 1949. It consisted of four columns: Scientific articles, Problems and exercises, Critics and bibliography and Meetings of the Society. Eight articles were published in this issue—we give their original titles.

1. J. Karamata: O aproksimaciji eksponencijalne funkcije nizom racionalnih funkcija

2. S. Fempl: O gornjoj granici razlike integrala proizvoda i proizvoda integrala 3. -D. Karapandžić: Prilog integraciji jedne obične diferencijalne jednačine

4. M. Tomić: Gauss-ov stav o težištu i njegova primena

5. I. Izbekov: O jednoj osobini prostih brojeva

6. D. Mihajlović: Partikularni integrali u problemu sudara triju tela, nebeske mehanike

7. D. Milosavljević: Skretanje tela ka istoku pri slobodnom padanju

8. D. Marković: O jednom istoriskom obrascu za kvadratni koren nekog broja

All the articles had abstracts in French and Russian.

In the column Meetings of the Society in this and subsequent issues (up to 1963), reports on all the important activities of the Society and, later, of the Union of Societies of Mathematicians, Physicists and Astronomers of Yugoslavia can be found. Thus, we find the report from the Founding session dated 04.01.1948, from the first session of the Society dated 26.04.1948, the third session dated 10.06.1948, dedicated to the fifth anniversary of the death of Mihailo Petrović, the detailed report from the First congress of the Union held on Bled 08–12.11.1949 and so on.

In the first year, two more issues of Vesnik were published—issue no. 2, as well as double-issue no. 3–4 for 1949. In the subsequent years, Vesnik continued to be published in single or double-issues. There were 147 issues in total, this issue no. 3–4 of the volume 50 for 1998 inclusive, the total number of pages being 16 474. The total number of scientific papers published in the main column Articles is 2017.

Starting from the issue 3–4 for 1950 the journal had a new editor-in-chief—it was Dragoljujb Marković, while the secreatry was Zlatko Mamuzić. The following issue, no. 1–2 for 1951 contained the first article written in foreign language; this was an article in French, written by Z. Mamuzić. In the subsequent years, the number of articles in foreign languages increased, and finally, starting from 1989, there were no more articles in Serbian. Besides, while in the beginning French was the most common language, later on there were more articles in Russian and German, and in the recent years about 90% of articles have been published in English.

In the first few years, only Yugoslav authors published in Vesnik. The first article with a foreign author can be found in the issue 3–4 for 1953. This was the article by E. Hille: Quelques remarques sur les equations de Kolmogoroff. The number of foreign authors increased later on, and so one can find a lot of very well known names among them; we mention only W. Sierpinski who published several articles in Vesnik and who was very active in the column Problems and exercises (later on Problems), too.

There was a reorganization, made after fifteen published volumes of Vesnik. By an agreement between the Society of Mathematicians, Physicists and Astronomers and the Mathematical Institute from Belgrade, the name of the journal changed to MATEMATIČKI VESNIK (“Mathematical bulletin”) and it began to be published by these two institutions. It had double numeration—the volume for 1964 was labeled as 1 (16). Dragoljub Marković continued to be editor-in-chief, and Milorad Bertolino was the secretary. The journal had the following six columns: Articles, Expositions of actual problems, Mathematical notes (till 1967), Problems, Book reviews and Scientific informations. Each column had its own editor.

After the death of Dragoljub Marković in the beginning of 1965, Zlatko Mamuzić was appointed the new editor-in-chief. Starting from 1968, Dušan Adna - dević was the secretary of the Editorial Board (substituted by Milorad Bertolino in 1971/72).

Some of the issues of Vesnik were dedicated to the articles from the scientific meetings, organized by the Society of Mathematicians or the Mathematical Institute. Thus, issue no. 4 (1968) contained the articles from the Symposium on Mihailo Petrović, on the occasion of his 100th anniversary; issues no. 1 (1985), 4 (1986) and 3–4 (1988) were dedicated to the Symposia on Complex Analysis, and issue no. 1 (1997) to the 11th Geometrical Symposium. Finally, issue no. 3–4 (1997) contained articles from the 4th Symposium on Mathematical Analysis and its Applications.

Another change of the publisher was made in 1977. Again, the Society of Mathematicians itself began to publish the journal, the editor-in-chief was still Zlatko Mamuzić, and the secretary was Vladimir Mićić. There were four columns: Articles and notes, Problems, Book reviews and Scientific informations. Till 1982 there was triple numeration of volumes (for example, 1 (14) (29) for 1977), and after that just single numeration showing the year of publishing (for example, vol. 35 for

The editor-in-chief was changed in 1979—the new editor was Dušan Adna - dević, and the secretaries were Zoran Kadelburg from 1982, and Pavle Mladenović from 1989. Finally, the last change of the editor was made in 1994—now the editorin-chief is Zoran Kadelburg, and Aleksandar Lipkovski is the secretary since the beginning of the present year.

The Editorial Board was refreshed several times. Starting from 1996, some foreign mathematicians were included in the Board, in an effort to raise the quality of articles. The present Editorial Board is:

Editor-in-chief : Zoran Kadelburg (Beograd)

Secretary: Aleksandar Lipkovski (Beograd)

Members:

Jan M. Aarts (Delft)  Petar Madić (Beograd)
Dušan Adamović (Beograd)  Vladimir Mićić (Beograd)
Dušan Adna - dević (Beograd)  Pavle Mladenović (Beograd)
Neda Bokan (Beograd)  Mila Mršević (Beograd)
Theodoros Bolis (Ioannina)  Predrag Obradović (Podgorica)
Dragan Vukotić (Madrid)  Miroslav Pavlović (Beograd)
Danica Despotović (Novi Sad)  Gordana Pavlović-Lažetić (Beograd)
Mirko Janc (Baltimore)  Alexander Šostak (Riga)
Boško Jovanović (Beograd)  Endre Süli (Oxford)
Petar Kenderov (Sofia)  Ioan Tomescu (Bucuresti)

From the very beginning Vesnik have been exchanged with a great number of journals all over the world. There are presently about hundred of such journals, including some of the most eminent ones. Due to lack of space, these journals are stored at the Faculty of Mathematics in Belgrade.

We can conclude that MATEMATIČKI VESNIK has played a very important role in the development of mathematical sciences in Yugoslavia. Some of the most eminent mathematicians published their articles in it, and, on the other hand, a lot of our mathematicians had an opportunity to publish their first articles in this journal. All the articles from Vesnik have been regularly reviewed in the main reviewing journals—Mathematical Reviews, Zentralblatt für Mathematik und ihre Grenzgebiete and ������������ ������. Finally, starting from 1996, Matematički vesnik is published electronically, too, as a part of the ELibEMS (Electronic Library of the European Mathematical Society) and it can be obtained through Internet on http://www.emis.de/journals/MV/ or http://www.matf.bg.ac.yu/dms/.

This is the last issue of the jubilar, 50th volume of Matematički Vesnik; we are sure that this jubilee will not be the last one.

Zoran Kadelburg